OpenStudy (anonymous):

Find all real solutions of the equation exactly 6z^4-13z^2+5=0

OpenStudy (anonymous):

6z^4-13z^2+5=0 6z^4-10z^2-3z^2+5=0 3z^2(2z^2-5)-1(2z^2-5)=0 (2z^2-5)((3z^2-1)=0 so either (2z^2-5)=0 or (3z^2-1)=0 z^2=5/2 or z^2=1/3 z=5/2 or z=-5/2 or z=1/3 or z=-1/3